OpenStudy (bgrg007):

how to solve for y' in (y')^2 + ty' + 4y=0. Please show the steps

OpenStudy (anonymous):

You have an equation that's quadratic in \(y'\). Complete the square: \[\begin{align*} (y')^2+ty'+4y&=0\\\\ (y')^2+ty'+\frac{t^2}{4}-\frac{t^2}{4}+4y&=0\\\\ \left(y'+\frac{t}{2}\right)^2-\frac{t^2}{4}+4y&=0\\\\ \end{align*}\]

OpenStudy (bgrg007):

clear explanation. thank you